Draw the curves for:

30 – Q = 2Q

3Q = 30 or Q = 10

at Q =10 P = 2*10 = 20
Answer #1

P = 2Q (This must be supply function as price and quantity have positive relation)

Thus Ps = 2Q

30 - Q = P (This must be demand function as price and quantity have inverse relation)

Thus Pd = 30 - Q

At equilibrium,

2Q = 30 - Q

Q = 10

P = 20

beinn 20 10 Quantity